The Big Bazinga Company makes molecular models. Budgeted costs to produce a model are as follows:

Input

Quantity

Price

Total Standard Cost

Direct materials

10 kg

$5.00 per kg

$ 50

Direct labour

5 hours

$25 per hour

125

Actual results for the year were as follows:

  • Produced 11,000 models
  • Direct material purchased:
    • 125,000 kg at $5.20/kg
  • Direct materials used:
    • 126,000 kg
  • Direct Labour: $1,622,500
  • Direct Labour Price (Rate) Variance is $147,500U

Required:

  1. Calculate the following:
    1. Direct Material Price (Rate) Variance
    2. Direct Material Quantity (Efficiency)Variance
    3. What was the amount of actual direct labour hours for the year?
    4. Direct Labour Quantity (Efficiency) Variance

  1. Give a reason for  the Direct Labour Quantity (Efficiency) Variance calculated above.
Answer #1
Req a)
i) Direct Material Price Variance = (Actual Price - Standard Price) * Actual Quantity
DMPV = (5.20 - 5) * 126000 = $25200 U
ii) Direct Material Quantity Variance = (Std. Quantity for actual production - actual quantity) * Std. price
DMQV = (11000*10 - 126000)*5 = $80000 U
iii) Direct Labor Rate Variance = (Actual rate - Standard rate) * Actual hours
Actual Labor - Actual hours * std. rate = DLRV
1622500 - Actual hours * 25 = 147500
Actual hours = (1622500 - 147500)/25 = 59000 hours
iv) Direct Material Efficiency Variance = (Std. hours for actual production - actual hours) * Std. rate
DMEV = (11000*5 - 59000) * 25 = $100000 U
Req b)
The efficiency of the labor is bad because untrained and inexperienced labors are
on the job which leads to good idle hours without being productive.